Mr. Castro walked from his house to the park at 2 kph and return by trisikle at 8 kph his return journey takes 2 and a half hour less than his outward journey. Find the number of hours he spent on the road?
Answer by Nate(3500)

Outside of town:
rate = 2
time = t
distance = 2t
Into town:
rate = 8
time = t - 2.5
distance = 8t - 20
The distances were equal ...
2t = 8t - 20
20 = 6t
10/3 = t
t - 2.5 = 5/6
TOTAL TIME = 10/3 + 5/6 = 20/6 + 5/6 = 25/6